OpenStudy (anonymous):

simplify: 2/3x-2= -3/2x-4 i definitely messed up somewhere solving this

hero (hero):

Is it: \(\large\frac{2}{3x-2} = -\frac{3}{2x-4}\)

OpenStudy (anonymous):

no it is 2/3x-2 = \[-3/2x-4\]

hero (hero):

That's the same thing as above, trust me.

hero (hero):

You should probably use the drawing tool next time to post your problem because what you just posted can be interpreted in different ways.

OpenStudy (anonymous):